OpenStudy (anonymous):

How many terms are in this series: 1 – 2 + 4 – 8 + . . . – 131072

OpenStudy (is3535):

-0

OpenStudy (perl):

this is a geomeric series

OpenStudy (perl):

geometric*

OpenStudy (perl):

if you divide two consecutive terms, you get a ratio of -2

OpenStudy (perl):

$$\Large (-2)^0 + (-2)^1 + (-2)^2 + (-2)^3+ ... (-2)^{n-1} $$

OpenStudy (anonymous):

an=a1(r)^n-1

OpenStudy (perl):

so you need to solve (-2)^(n-1) = – 131072

OpenStudy (perl):

you can also solve this using logarithms

OpenStudy (anonymous):

n=18

OpenStudy (anonymous):

Determine the sum: 25 + 5 + 1 + ...

OpenStudy (perl):

the ratio is 1/5

OpenStudy (perl):

$$\Large \sum_{n=0}^{\infty} ar^{n-1}= \frac{a}{1-r}~,~ |r| < 1 $$

OpenStudy (anonymous):

a1=25

OpenStudy (perl):

right

OpenStudy (anonymous):

25/1/5

OpenStudy (perl):

be careful

OpenStudy (perl):

$$\Large \sum_{n=0}^{\infty} ar^{n-1}= \frac{a}{1-r}~,~ |r| < 1 \\~\\\Large \sum_{n=0}^{\infty} 25\cdot (1/5)^{n-1}= \frac{25}{1-1/5}~,~ |1/5| < 1 ~\checkmark \\~\\ \Large \frac{25}{4/5} = 25 \cdot \frac 5 4 $$

OpenStudy (anonymous):

31.25!

OpenStudy (anonymous):

i almost forgot to do 1-1/5

OpenStudy (anonymous):